OpenStudy (calculusxy):

If A + B ------- = 4 3 and A is greater than 1, which of the following could NOT be the value of B? A. -3 B. 0 C. 1 D. 2 E. 12 @ganeshie8

OpenStudy (calculusxy):

Is it A.

ganeshie8 (ganeshie8):

lets see :) A + B ------- = 4 3 solve B

ganeshie8 (ganeshie8):

A + B ------- = 4 3 A + B = 12 B = 12-A

ganeshie8 (ganeshie8):

since A is > 1, B < 12 - 1 B < 11

OpenStudy (calculusxy):

Please I am in a hurry

ganeshie8 (ganeshie8):

you're also looking for ansurs ha ;) answer is E. 12

OpenStudy (charlotte123):

-3 - 0 - 1 ^ ^ ^ These numbers are not greater than one. And 2/3 ≠ 4

OpenStudy (charlotte123):

Thus - A Is not the answer. Clearly the answer Is not B or C. And D Is wrong as well................